What is the 97.5th Percentile in a t distribution with a sample size of 15?

1
Expert's answer
2022-05-05T07:28:20-0400

We use the t-table, the degrees of freedom df = n - 1 = 15 - 1 =14,

the 97.5th percentile is 2.145.